if sin(theta)=3/5 and (theta) is a second quadrant angle, find tan2(theta)
Answer by Alan3354(21606)

if sin(theta)=3/5 and (theta) is a second quadrant angle, find tan2(theta)
-----------
cos(theta) = sqrt(1 - sin(theta))
cos(theta) = 4/5, -4/5
tan = (3/5)/(4/5) = 0.75 or -0.75
Since it's 2nd quadrant, it's -0.75
------------------------
I don't know if the 2 is an exponent or a multiplier.
You'll have to finish it.